OpenStudy (anonymous):

please help step by step...Identify the correct slope and y intercept of the equation 6x - 3y = 12. slope = 2; y intercept at (0, -4) slope = -2; y intercept at (0, 4) slope = one half; y intercept at (0, 15) slope = one half; y intercept at (0, 9)

OpenStudy (the_fizicx99):

6x - 3y = 12 -6x -6x -3y = -6x + 12 /-3 /-3 /-3 y = 6/3x -4 Thats y = mx + b B is the y intercept

OpenStudy (the_fizicx99):

6/3 = 2x

OpenStudy (anonymous):



OpenStudy (the_fizicx99):

So A

OpenStudy (anonymous):

oh i remember now I just need a little refresher

OpenStudy (the_fizicx99):

@dalia_lam you didn't exlpain how to get Ax + By = c to y = mx + b

OpenStudy (anonymous):

its fine you did

OpenStudy (anonymous):